Train A and B are traveling parallel. Train a is traveling 100 miles per hour and tain b is traveling at 125 miles per hour. train a passes the statin at 7:10 pm and train b passes at 7:22 pm. waht time will train b cath up to train a?
Answer by checkley77(12844)

D=RT
D=100T for train A.
D=125(T-.2) for train B.
100T=125(T-.2)
100T=125T-25
100T-125T=-25
-25T=25
T=25/25
T=1 HOUR ANS.
PROOF:
100*1=125(1-.2)
100=125*.8
100=100